Session: Oct 2017 (session 3)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (C.Bortolotti at Sun Oct 29 15:06:48 2017) Success:
Observation OK

Yebes (Daniel Valero at Mon Oct 30 08:16:11 2017) Success:
Not known problem

Hartebeesthoek (Jonathan Quick at Mon Oct 30 08:19:09 2017) Success:
No known problems.

Onsala (Jun Yang at Mon Oct 30 13:35:58 2017) Minor failures:
This NME observations and the next scientific user experiment requested to run the observations at about 21.7 GHz. With the default LO=21.5 GHz, this observing frequency is outside of the valid bandwidth (IF filter: 500 - 1000 MHz). The problem was found after the first ftp-fringe test. To get the right observing window, a new LO value of 21 GHz was used and the schedule was restarted. The telescope was on source again at 13:43:37 UT. After that, fringes to Onsala were found.

Effelsberg (Uwe at Thu Nov 02 11:07:43 2017) Minor failures: WIND
Started 45 min late, because of strong wind.